OpenStudy (anonymous):

Solve the following system of equations. 2x – 3y + z = –1 3x + 2y + 2z = –1 x – y – 3z = –4

OpenStudy (texaschic101):

2x - 3y + z = -1 -->(-2) 3x + 2y + 2z = -1 -------------------- -4x + 6y - 2z = 2 (result of multiplying by -2) 3x + 2y + 2z = -1 --------------------add -x + 8y = 1 3x + 2y + 2z = -1 -->(3) x - y - 3z = -4 -->(2) -------------------- 9x + 6y + 6z = -3 (result of multiplying by 3) 2x - 2y - 6z = -8 (result of multiplying by 2) --------------------add 11x + 4y = -11 -x + 8y = 1 11x + 4y = -11 -->(-2) --------------------- -x + 8y = 1 -22x - 8y = 22 (result of multiplying by -2) ------------------- -23x = 23 x = -1 -x + 8y = 1 -(-1) + 8y = 1 1 + 8y = 1 8y = 1-1 8y = 0 y = 0/8 y = 0 2x - 3y + z = -1 2(-1) - 3(0) + z = -1 -2 - 0 + z = -1 -2 + z = -1 z = -1 + 2 z = 1 check..... 3x + 2y + 2z = -1 3(-1) + 2(0) + 2(1) = -1 -3 + 0 + 2 = -1 -3 + 2 = -1 -1 = -1 (correct) x = -1, y = 0, z = 1 or (-1,0,1)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

holy crap you are really good at this thank you so much!

OpenStudy (texaschic101):

you are very welcome...glad I could help :)